a uniform rod AB of length a and weight w is freely hinged to a vertical wall at A and is maintained in equlibrium by a light string of lenght a fastened to B and to a point C at a distance b vertically above A. prove reaction at hinge WSqrt(a2-2b2)/2b and find tension in the string


1
Expert's answer
2021-01-27T10:17:33-0500

Solution:


About the hinge, the torque due to the tension in the string is equal to the torque due to the weight:

T*sin(2"\theta" )*a = W*("\tfrac{a}{2}")*cos"\theta"

substitute sin(2"\theta" ) = 2sinΘcosΘ and cancel the a*cos"\theta" terms:

T*2*sin"\theta" = "\dfrac{W}{2}"

T = "\dfrac{W}{4sin\theta}"


But sin"\theta" = "\tfrac{b}{2a}" , so


T = "\dfrac{W}{4(\frac{b}{2a})}=\dfrac{aW}{2b}"

which has vertical component

Ty = T*sinΘ = "(\dfrac{aW}{2b})*(\dfrac{b}{2a})=\dfrac{W}{4}"

and horizontal component

Tx = T*cosΘ = "(\dfrac{aW}{2b})"*"\dfrac{ \sqrt{\smash[b]{(a² - \frac{b^2}{4})}}}{a}" = "(\dfrac{W}{2b})" *"\sqrt{\smash[b]{(a² - \frac{b^2}{4})}}"

That makes the reaction components at the hinge

Rx = Tx

and Ry = W - "\dfrac{W}{4}" = "\dfrac{3W}{4}".

Then

R = "\\sqrt{\\smash[b]{[(\\dfrac{W}{2b})\n\u200b\t\n \u00b2*(a\u00b2 - \\frac{b^2}{4} \n\n \n\u200b\t\n ) + (\\frac{3W}{4} \n\u200b\t\n )\u00b2]}}"


R = "(\dfrac{W}{2b})" *"\\sqrt{\\smash[b]{[a\u00b2 - \\frac{b^2}{4} \n\n \n\u200b\t\n + \\frac{9b^2}{4} \n\n\u200b\t\n ]}}"


R = "(\dfrac{W}{2b})" *"\sqrt{\smash[b]{[a² + 2b²]}}"